OpenStudy (anonymous):

Determine the nature of solutions of the equation x^2-3x+8=0

OpenStudy (anonymous):

you need to check the discriminant - know what that is ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

are you familiar with the quadratic equation ?

OpenStudy (anonymous):

use quadratic formula

OpenStudy (anonymous):

no not at all.

OpenStudy (anonymous):

ever seen this formula ? \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^{2}-4ac})/2a\]

OpenStudy (anonymous):

In general: ax^2+bx+c=0 In your case: a=1 (the coefficient of x^2) b=-3 (coefficient of x) c=8 Use these values and plug them into the discriminant (the thingy inside the square root). \[\Delta=b^2-4ac\] If the discriminant = 0, there is one real solution If the discriminant > 0, there are two real solutions If the discriminant < 0, there are no real solutions